OpenStudy (anonymous):

Jessica is 3 years older than Rebecca. In 5 years, she will be 3 times as old as Rebecca was 2 years ago. Find the girls' present ages.

Parth (parthkohli):

Let's try it with an equation. If Rebecca is \(x\), then Jessica is \(x + 3\) right?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Now, just make an equation that looks like: \(x + 3 + 5 = 3(x - 2)\) Got it?

Parth (parthkohli):

Then solve the equation. \( \color{Black}{\Rightarrow x + 8 = 3(x - 2)}\) \( \color{Black}{\Rightarrow x + 8 = 3x - 6}\) Continue, please.

OpenStudy (ganpat):

let present age be of Rebecca be x SO, Jessica = x + 3 (x+3) + 5 = 3 (x -2) x+ 3+ 5 = 3x - 6 x + 8 = 3x - 6 2x = 14 x = 7 rebecca = 7 and Jessica = 7 + 3 = 10...

OpenStudy (ganpat):

@ParthKohli : shd not it be (x+3) + 5 = 3 (x -2) + 5

Parth (parthkohli):

I am sorry, but Ganpat is right :)

Parth (parthkohli):

\(x + 3 + 5 = 3(x - 2) + 5\)

OpenStudy (anonymous):

Oh I understand now! Thanks guys @Ganpat @ParthKohli

OpenStudy (ganpat):

no. i also went wrong :-p , i realized later :D

OpenStudy (ganpat):

W/c !!